# Jaynes-Cummings model in a unitary fractional-time description

### v1

@@ II.3 Unitary fractional-time evolution
where the Dyson map acting on the two-dimensional subspace $\mathcal{H}_{n}$ is chosen as being
@@
where we assume that $\lambda_{\alpha}^{(n)}(t)\in\mathbb{C}$ and $\kappa_{\alpha}^{(n)}(t),\Lambda_{\alpha}^{(n)}(t)\in\mathbb{R}$ under the additional condition $\Lambda_{\alpha}^{(n)}(t)>0$. Moreover, we can also represent the $n$-th subspace of Dyson map as a matrix in the basis $\{|e,n\rangle,|g,n+1\rangle\}$, it yields
$\hat{\eta}_{\alpha}^{(n)}(t)=\frac{e^{\kappa_{\alpha}^{(n)}}}{\sqrt{\Lambda_{\alpha}^{(n)}}}\left[\begin{matrix}\chi_{\alpha}^{n}&\lambda_{\alpha}^{(n)}\\ [\lambda_{\alpha}^{(n)}]^{\ast}&1\end{matrix}\right],$ (16)
where we omit the time dependencies in the Dyson map parameters to clean up the notation and define the function $\chi_{\alpha}^{(n)}(t)=\Lambda_{\alpha}^{(n)}+|\lambda_{\alpha}^{(n)}|^{2}$.
In applying the results of Eqs. (11) and (16) into Eq. (14), we have the operator $\hat{u}_{\alpha}(t)$ to be represented as
@@
with $\hat{u}_{\alpha}^{(n)}(t)=\hat{\eta}_{\alpha}^{(n)}(t)\hat{U}_{\alpha}^{(n)}(t)[\hat{\eta}_{\alpha}^{(n)}(0)]^{-1}$. Admitting $\hat{u}_{\alpha}(t)$ as a unitary operator, implies on $\hat{u}_{\alpha}^{(n)}(t)$ must necessarily belong to the Lie group $U(2)$, and it can be represented in the matrix form
$\hat{u}^{(n)}_{\alpha}(t)=e^{i\delta_{\alpha}^{(n)}}\left[\begin{matrix}\varpi_{\alpha,+}^{(n)}&\varpi_{\alpha,-}^{(n)}\\ -[\varpi_{\alpha,-}^{(n)}]^{\ast}&[\varpi_{\alpha,+}^{(n)}]^{\ast}\end{matrix}\right],$ (18)
@@
$\varpi_{\alpha,\pm}^{(n)}(t)={}$$\pm e^{i\delta_{\alpha}^{(n)}}[\nu_{\alpha,\mp}^{(n)}]^{\ast},$ (19a)
$\delta_{\alpha}^{(n)}(t)={}$ $\frac{1}{2}\mathrm{Im}[\ln{D_{\alpha}^{(n)}}],$ (19b)
in which $D_{\alpha}^{(n)}(t)=[\mathcal{C}^{(n)}_{\alpha}]^{2}-(-1)^{-\alpha}[\mathcal{S}^{(n)}_{\alpha}]^{2}$, and the coefficients satisfying the relation $|\varpi_{\alpha,+}^{(n)}|^{2}+|\varpi_{\alpha,-}^{(n)}|^{2}=1.$ In Eq. (19), the functions $\nu_{\alpha,\pm}^{(n)}$ are given by
@@
$\nu_{\alpha,\pm}^{(n)}(t)={}$$\pm\frac{e^{\kappa_{\alpha}^{(n)}-\kappa_{\alpha}^{(n)}(0)}}{\sqrt{\Lambda_{\alpha}^{(n)}\Lambda_{\alpha}^{(n)}(0)}}\left[\zeta_{\alpha,\pm}^{(n)}+[\lambda_{\alpha}^{(n)}]^{\ast}\xi_{\alpha,\pm}^{(n)}\right],$ (20)
with the functions $\zeta_{\alpha,\pm}^{(n)}$ e $\xi_{\alpha,\pm}^{(n)}$ defined as
$\zeta_{\alpha,+}^{(n)}(t)$ $=i^{-\alpha}\mathcal{S}_{\alpha}^{(n)}-[\lambda_{\alpha}^{(n)}(0)]^{\ast}\mathcal{C}_{\alpha}^{(n)},$ (21a)
$\zeta_{\alpha,-}^{(n)}(t)$ $=i^{-\alpha}\lambda_{\alpha}^{(n)}(0)\mathcal{S}_{\alpha}^{(n)}-\chi_{\alpha}^{(n)}(0)\mathcal{C}_{\alpha}^{(n)},$ (21b)
$\xi_{\alpha,+}^{(n)}(t)$ $=\mathcal{C}_{\alpha}^{(n)}-i^{-\alpha}[\lambda_{\alpha}^{(n)}(0)]^{\ast}\mathcal{S}_{\alpha}^{(n)},$ (21c)
$\xi_{\alpha,-}^{(n)}(t)$ $=\lambda_{\alpha}^{(n)}(0)\mathcal{C}_{\alpha}^{(n)}-i^{-\alpha}\chi_{\alpha}^{(n)}(0)\mathcal{S}_{\alpha}^{(n)}.$ (21d)
From the imposition of unitarity of $\hat{u}_{\alpha}^{(n)}$, the Dyson map parameters have the form
$\kappa_{\alpha}^{(n)}(t)={}$ $\kappa_{\alpha}^{(n)}(0)-\frac{1}{2}\mathrm{Re}[\ln{D_{\alpha}^{(n)}}],$ (22a)
$\chi_{\alpha}^{(n)}(t)={}$ $\frac{|\zeta_{\alpha,+}^{(n)}|^{2}+|\zeta_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}}{|\xi_{\alpha,+}^{(n)}|^{2}+|\xi_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}},$ (22b)
$\lambda_{\alpha}^{(n)}(t)={}$ $\frac{-\left[\xi_{\alpha,+}^{(n)}[\zeta_{\alpha,+}^{(n)}]^{\ast}+\xi_{\alpha,-}^{(n)}[\zeta_{\alpha,-}^{(n)}]^{\ast}\right]}{|\xi_{\alpha,+}^{(n)}|^{2}+|\xi_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}},$ (22c)
@@
where the time-dependent Dyson map parameter $\Lambda_{\alpha}^{(n)}=\chi_{\alpha}^{(n)}-|\lambda_{\alpha}^{(n)}|^{2}$, in which we must have $\chi_{\alpha}^{(n)}>|\lambda_{\alpha}^{(n)}|^{2}$, once we are assuming $\Lambda_{\alpha}^{(n)}$ as a positive function. Notice that these functions depend only on the fractional-time evolution parameters and on the initial values of Dyson map parameters into the functions $\xi_{\alpha,\pm}^{(n)}(t)$ and $\zeta_{\alpha,\pm}^{(n)}(t)$. For more details about those calculations see Ref. [37].
III Atomic Population Inversion: collapse and revival
@@
The phenomenon of collapse and revival of atomic oscillations is a distinctive feature observed in the interaction of a two-level atom with a quantized electromagnetic field inside a cavity described by the JC model. When the cavity field is prepared in a coherent state and interacts with the two-level atom, the system undergoes a fascinating dynamical behavior characterized by the following stages: i) initial Rabi Oscillations: at the outset, the atom exchanges energy with the quantized field, resulting in oscillations of the atomic population between the ground and excited states. These are known as Rabi oscillations; ii) collapse: due to the quantized nature of the field and the distribution of photon number states in the coherent state, these oscillations begin to dephase. This dephasing causes the observable oscillations to diminish in amplitude, leading to a phenomenon known as the "collapse" of the Rabi oscillations; iii) revival: after a certain period, the oscillations rephase, and the atomic population oscillations reappear. This re-emergence of oscillations is termed the "revival" of the Rabi oscillations. The timing of these revivals is determined by the properties of the coherent state and the parameters of the system.
@@
We aim to analyze this phenomenon depending on the $\alpha$ parameter in the unitary approach. In what follows, we consider the situation in which the atom is initially prepared in the excited state and the field is a coherent state, such that $|\psi_{\alpha}(0)\rangle=|e\rangle\otimes|\beta\rangle$, where $|\beta\rangle=\sum_{n=0}^{\infty}c_{n}|n\rangle$ being $c_{n}=e^{-|\beta|^{2}/2}\beta^{n}/\sqrt{n!}$ with $\beta\in\mathbb{C}$. The system evolves in according to $|\psi_{\alpha}(t)\rangle=\hat{u}_{\alpha}(t)|\psi_{\alpha}(0)\rangle$ that results in
@@
where the time-dependent probability amplitudes are
@@
The probability of finding the atom in the excited state with the field having $n$ photons is $P_{e,n}^{\alpha}(t)=|A_{e,n}^{\alpha}(t)|^{2}$. In contrast, the probability of finding the atom in the ground state with the field having $n+1$ photons is $P_{g,n+1}^{\alpha}(t)=|A_{g,n}^{\alpha}(t)|^{2}$. Furthermore, we can marginalize the probability over the field states, by summing over all possible photon numbers to obtain the probability of finding the atom in an excited or grounded state, which are given respectively by $P_{e}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{e,n}^{\alpha}(t)|^{2}$ and $P_{g}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{g,n}^{\alpha}(t)|^{2}$ We then calculate the population inversion of the atom along the time, which is given by the mean value of the inversion operator,
@@
Fig. 1 illustrates the results for the coherent state parameter $\beta=2$ and various values of the parameter $\alpha$, with initial conditions for the Dyson map set as $\kappa_{\alpha}^{(n)}(0)=0$, $\Lambda_{\alpha}^{(n)}(0)=1$, and $\lambda_{\alpha}^{(n)}(0)=0$. These conditions correspond to the scenario where $\hat{\eta}_{\alpha}(0)$ is the identity operator, ensuring that the initial states in both the Hermitian and non-Hermitian representations are identical, i.e., $|\psi_{\alpha}(0)\rangle=|\Psi_{\alpha}(0)\rangle$. For $\alpha=1.00$ (black solid line), we observe the well-known result of population inversion dynamics, as the fractional derivative reduces to the standard first-order derivative in Eq. (1). When $\alpha=0.75$ (blue dashed line), the population inversion exhibits a slightly modified behavior compared to the usual case ($\alpha=1.00$), though with fewer oscillations. In contrast, for $\alpha=0.50$ (orange dotted line) the rapid oscillations in population inversion vanish, giving rise to a more periodic behavior. Additionally, the population eventually returns to the excited state, indicating that the combination of the fractional-derivative parameter and the Dyson map introduces additional driving terms in the atom-field interaction within the unitary description.
Figure 1: Time-evolution of the atomic population inversion of the atom for different values of the fractional-order parameter $\alpha$. We plot the usual case corresponding to $\alpha=1.0$ (black solid line), $\alpha=0.75$ (blue dashed line) and $\alpha=0.5$ (orange dotted line). We start with the atom in the excited state and the field in the coherent state with $\beta=2$, and the initial values of Dyson map parameters being $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$.
IV Entanglement between atom-field

### v2

@@ III.3 Unitary evolution
with the TD Dyson map acting on the two-dimensional subspace $\mathcal{H}_{n}$ chosen as
@@
where we assume $\lambda_{\alpha}^{(n)}(t)\in\mathbb{C}$ and $\kappa_{\alpha}^{(n)}(t),\Lambda_{\alpha}^{(n)}(t)\in\mathbb{R}$ with the additional condition $\Lambda_{\alpha}^{(n)}(t)>0$. Furthermore, we can represent the $n$-th subspace of the time-dependent Dyson map as a $2\times 2$ matrix in the basis $\{|e,n\rangle,|g,n+1\rangle\}$, yielding:
$\hat{\eta}_{\alpha}^{(n)}(t)=\frac{e^{\kappa_{\alpha}^{(n)}(t)}}{\sqrt{\Lambda_{\alpha}^{(n)}(t)}}\left[\begin{matrix}\chi_{\alpha}^{n}(t)&\lambda_{\alpha}^{(n)}(t)\\ [\lambda_{\alpha}^{(n)}(t)]^{\ast}&1\end{matrix}\right]{\color[rgb]{1,0,0}{,}}$ , (19)
being the function $\chi_{\alpha}^{(n)}(t)=\Lambda_{\alpha}^{(n)}(t)+|\lambda_{\alpha}^{(n)}(t)|^{2}$.
By applying the results of Eqs. (16) and (19) in Eq. (7), we represent the unitary operator $\hat{u}_{\alpha}(t)$ as
@@
with $\hat{u}_{\alpha}^{(n)}(t)=\hat{\eta}_{\alpha}^{(n)}(t)\hat{U}_{\alpha}^{(n)}(t)[\hat{\eta}_{\alpha}^{(n)}(0)]^{-1}$. Admitting $\hat{u}_{\alpha}(t)$ as a unitary operator, it implies that $\hat{u}_{\alpha}^{(n)}(t)$ must necessarily belong to the Lie group $U(2)$ and can be represented in the general matrix form
$\hat{u}^{(n)}_{\alpha}(t)=e^{i\delta_{\alpha}^{(n)}(t)}\left[\begin{matrix}\varpi_{\alpha,+}^{(n)}(t)&\varpi_{\alpha,-}^{(n)}(t)\\ -[\varpi_{\alpha,-}^{(n)}(t)]^{\ast}&[\varpi_{\alpha,+}^{(n)}(t)]^{\ast}\end{matrix}\right],$ (21)
@@
$\delta_{\alpha}^{(n)}(t)={}$ $\frac{1}{2}\mathrm{Im}[\ln{D_{\alpha}^{(n)}(t)}],$ (22a)
$\varpi_{\alpha,\pm}^{(n)}(t)={}$$\pm e^{i\delta_{\alpha}^{(n)}(t)}[\nu_{\alpha,\mp}^{(n)}(t)]^{\ast},$ (22b)
where the function $D_{\alpha}^{(n)}(t)$ in Eq. (22a) is defined as $D_{\alpha}^{(n)}(t)=[\mathcal{C}^{(n)}_{\alpha}(t)]^{2}-(-1)^{-\alpha}[\mathcal{S}^{(n)}_{\alpha}(t)]^{2}$. To simplify the notation, from now on we omit the time dependence and define $\nu_{\alpha,\pm}^{(n)}$ in Eq. (22b) as follows:
@@
$\nu_{\alpha,\pm}^{(n)}={}$$\pm\frac{e^{\kappa_{\alpha}^{(n)}-\kappa_{\alpha}^{(n)}(0)}}{\sqrt{\Lambda_{\alpha}^{(n)}\Lambda_{\alpha}^{(n)}(0)}}\left[\zeta_{\alpha,\pm}^{(n)}+(\lambda_{\alpha}^{(n)})^{\ast}\xi_{\alpha,\pm}^{(n)}\right],$ (23)
where $\zeta_{\alpha,\pm}^{(n)}$ and $\xi_{\alpha,\pm}^{(n)}$ are given by
$\zeta_{\alpha,+}^{(n)}$ $=i^{-\alpha}\mathcal{S}_{\alpha}^{(n)}-[\lambda_{\alpha}^{(n)}(0)]^{\ast}\mathcal{C}_{\alpha}^{(n)},$ (24a)
$\zeta_{\alpha,-}^{(n)}$ $=i^{-\alpha}\lambda_{\alpha}^{(n)}(0)\mathcal{S}_{\alpha}^{(n)}-\chi_{\alpha}^{(n)}(0)\mathcal{C}_{\alpha}^{(n)},$ (24b)
$\xi_{\alpha,+}^{(n)}$ $=\mathcal{C}_{\alpha}^{(n)}-i^{-\alpha}[\lambda_{\alpha}^{(n)}(0)]^{\ast}\mathcal{S}_{\alpha}^{(n)},$ (24c)
$\xi_{\alpha,-}^{(n)}$ $=\lambda_{\alpha}^{(n)}(0)\mathcal{C}_{\alpha}^{(n)}-i^{-\alpha}\chi_{\alpha}^{(n)}(0)\mathcal{S}_{\alpha}^{(n)},$ (24d)
Moreover, for $\hat{u}_{\alpha}^{(n)}$ to represent a unitary operator, the parameters of the TD Dyson map must take the form:
$\kappa_{\alpha}^{(n)}={}$ $\kappa_{\alpha}^{(n)}(0)-\frac{1}{2}\mathrm{Re}[\ln{D_{\alpha}^{(n)}}],$ (25a)
$\chi_{\alpha}^{(n)}={}$ $\frac{|\zeta_{\alpha,+}^{(n)}|^{2}+|\zeta_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}}{|\xi_{\alpha,+}^{(n)}|^{2}+|\xi_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}},$ (25b)
$\lambda_{\alpha}^{(n)}={}$ $\frac{-\left[\xi_{\alpha,+}^{(n)}[\zeta_{\alpha,+}^{(n)}]^{\ast}+\xi_{\alpha,-}^{(n)}[\zeta_{\alpha,-}^{(n)}]^{\ast}\right]}{|\xi_{\alpha,+}^{(n)}|^{2}+|\xi_{\alpha,-}^{(n)}|^{2}+\Lambda_{\alpha}^{(n)}(0)e^{\mathrm{Re}[\ln{D_{\alpha}^{(n)}}]}},$ (25c)
@@
where the TD Dyson map parameter $\Lambda_{\alpha}^{(n)}=\chi_{\alpha}^{(n)}-|\lambda_{\alpha}^{(n)}|^{2}$, with the condition that $\chi_{\alpha}^{(n)}>|\lambda_{\alpha}^{(n)}|^{2}$, since we are assuming $\Lambda_{\alpha}^{(n)}$ as a positive function. The functions in Eq. (22b) satisfy the relation $|\varpi_{\alpha,+}^{(n)}|^{2}+|\varpi_{\alpha,-}^{(n)}|^{2}=1$, which comes from the fact of $|\det\hat{u}^{(n)}_{\alpha}|^{2}=1$. Note that the coefficients of the unitary operator $\hat{u}^{(n)}_{\alpha}$ depend only on the parameters of the fractional-time evolution operator and the initial values of the parameters of the TD Dyson map, which are incorporated into the functions $\xi_{\alpha,\pm}^{(n)}$ and $\zeta_{\alpha,\pm}^{(n)}$. For more details on those calculations, see Ref. [32].
IV Population Inversion
Figure 1: Time evolution of the population inversion $W_{\alpha}(t)$ for different values of $\alpha$. It is considered initially the atom in the excited state and the field in the coherent state with $\beta=2$, and the initial values of Dyson map parameters being $\kappa_{\alpha}(0)=0$, $\Lambda_{\alpha}(0)=1$ and $\lambda_{\alpha}(0)=0$. The atom-field coupling is $\mu_{\alpha}=1$. Then, we analyze for $\alpha=1.00$ (a), $\alpha=0.75$ (b), $\alpha=0.60$ (c), and $\alpha=0.50$ (d), $\alpha=0.45$ (e), $\alpha=0.40$ (f), $\alpha=0.30$ (g), and $\alpha=0.25$ (h).
@@ IV Population Inversion
The collapse and revival of atomic oscillations is a distinctive feature observed in the interaction of a two-level atom with a quantized electromagnetic field inside a cavity described by the JC model [46]. When the cavity field is initially prepared in a coherent state and interacts with the two-level atom, the system exhibits a fascinating dynamical behavior characterized by the following stages: i) initial Rabi oscillations: at the beginning, the atom exchanges energy with the quantized field, resulting in oscillations of the atomic population between the ground and excited states; ii) collapse: the distribution of photon number states within the coherent field leads to dephasing, causing the Rabi oscillations to decay; iii) revival: after a certain period, the oscillations rephase, and the atomic population oscillations reappear. The timing of these revivals is determined by the properties of the coherent state and the parameters of the system.
@@
We aim to analyze the atomic population inversion depending on the $\alpha$ parameter within a unitary framework. We consider an initial state where the atom is excited and the field is in a coherent state: $|\psi_{\alpha}(0)\rangle=|e\rangle\otimes|\beta\rangle$, where $|\beta\rangle=\sum_{n=0}^{\infty}c_{n}|n\rangle$ where $c_{n}=e^{-|\beta|^{2}/2}\beta^{n}/\sqrt{n!}$ with $\beta\in\mathbb{C}$. The system evolves in according to $|\psi_{\alpha}(t)\rangle=\hat{u}_{\alpha}(t)|\psi_{\alpha}(0)\rangle$, leading to
@@
where the probability amplitudes are
@@
The probability of finding the atom in the excited state with the field having $n$ photons is $P_{e,n}^{\alpha}(t)=|A_{e,n}^{\alpha}(t)|^{2}$. In contrast, the probability of finding the atom in the ground state with the field having $n+1$ photons is $P_{g,n+1}^{\alpha}(t)=|A_{g,n}^{\alpha}(t)|^{2}$. Furthermore, we can marginalize the probability over the field states, by summing over all possible photon numbers to obtain the probability of finding the atom in an excited or grounded state, which are given respectively by $P_{e}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{e,n}^{\alpha}(t)|^{2}$ and $P_{g}^{\alpha}(t)=\sum_{n=0}^{\infty}|A_{g,n}^{\alpha}(t)|^{2}$ We then calculate the population inversion over time, which is given by the mean value of the inversion operator,
@@
In Fig. 1, we illustrate the time evolution of the population inversion $W_{\alpha}(t)$ by considering the atom-field coupling $\mu_{\alpha}=1$ for a coherent state with $\beta=2$ and different values of the parameter $\alpha$. The initial conditions for the TD Dyson map are set as $\kappa_{\alpha}^{(n)}(0)=0$, $\Lambda_{\alpha}^{(n)}(0)=1$, and $\lambda_{\alpha}^{(n)}(0)=0$. These conditions imply that the Dyson map at $t=0$, $\hat{\eta}_{\alpha}(0)$, is the identity operator. In this case, the initial states in both the unitary and nonunitary representations are identical, thus, $|\psi_{\alpha}(0)\rangle=|\Psi_{\alpha}(0)\rangle$. As $\alpha$ decreases, the dynamics begin to change. For instance, when $\alpha=1.00$ (Fig. 1a), we recover the well-known population inversion dynamics of the JC model, as the Caputo derivative reduces to the first-order derivative in Eq. (1). For $\alpha=0.75$ (Fig. 1b), the population inversion exhibits a slightly modified behavior with a longer collapse time and fewer oscillations. For $\alpha=0.60$ (Fig. 1c), irregular collapse and revival patterns emerge. Interestingly, when $\alpha=0.50$ (Fig. 1d), the population inversion exhibits a more periodic behavior, returning close to the excited state with a period of approximately $6.18$. One might expect that further decreasing $\alpha$ would lead to even more regular behavior. However, for $\alpha<0.5$ the dynamics become highly irregular. Sudden peaks or troughs characterize the population inversion, and the characteristic collapse and revival patterns diminish due to the large fluctuations. For instance, when $\alpha=0.45$ (Fig. 1e), the regular pattern observed for $\alpha=0.50$ disappears. A near-complete population inversion occurs after the initial collapse, followed by fewer subsequent collapses. For $\alpha=0.40$ (Fig. 1f) and $\alpha=0.30$ (Fig. 1f) an irregular oscillations pattern with faster oscillations manifests. When $\alpha=0.25$, the population inversion exhibits a pattern of faster oscillations, but the overall dynamics remain quasi-periodic.
Remarkably, within the unitary framework, the unique properties of fractional derivatives can be attributed to additional driving forces in the atom-field interaction [51]. These forces arise from combining the fractional-derivative parameter and the TD Dyson map. This can be inferred by representing the system in the $\mathcal{H}_{n}$ basis, where the Hermitian Hamiltonian in Eq. (9) takes the form $\hat{h}_{\alpha}^{(n)}(t)\propto\mu_{i,\alpha}^{(n)}(t)\hat{\sigma}_{i}$ with $i=x,y,z$.
V Atom-field Entanglement